OpenStudy (0_0youscareme):

find the solution of 5b-13=22 if the replacement set is {5, 6, 7, 8, 9}

OpenStudy (0_0youscareme):

@pooja195

OpenStudy (anonymous):

7

OpenStudy (0_0youscareme):

can you explain how you get 7

OpenStudy (anonymous):

add 13 to both sides: 5b-13+13=22+13 5b=35 divid both sides by 5 b=7

OpenStudy (anonymous):

reply if helped :)

OpenStudy (0_0youscareme):

yes tyvm

OpenStudy (anonymous):

yw :)